A pile of 15 notebooks is111/2cm high. What is the thickness of each notebook?
Answer by stanbon(75887)

A relief organization supplies blankets, cots, and lanterns to victims of fires, floods, and other natural disasters. One week, the organization purchased 15 blankets, 5 cots, and 10 lanterns for a total cost of $1150. The next week, at the same prices, the organization purchased 20 blankets, 10 cots, and 15 lanterns for a total cost of $1800. The third week, at the same prices, the organization purchased 10 blankets, 15 cots, and 5 lanterns for a total cost of $1325. Find the cost of one blanket, the cost of one cot, and the cost of one lantern.
---------
Equations:
15b + 5c + 10L = 1150
20b + 10c + 15L = 1800
10b + 15c + 5L = 1325
-----------------------------
I used a matrix function on my TI-84 to get:
b = 25 (# of blankets)
c = 55 (# of cots)
L = 50 (# of lanterns)
=============================
Cheers,
Stan H.
===================